1) Calculate the momentum of the Titanic (m = 4.2 * 10^7 kg) moving at 14 knots (1 kt = 1.852 km/h)
Answer #1

here,

the mass of Titanic , m = 4.2 * 10^7 kg

the speed of Titanic , v = 1.852 km/h

v = 0.5144 m/s

the momentum of Titanic , P = m * v

P = 4.2 * 10^7 * 0.5144 kg.m/s

P = 2.16 * 10^7 kg.m/s